# Exploring Holomorphic Retracts

### v6

@@ 2.5. Fixed point sets and Retracts
If $D_{1},D_{2}$ are any pair of domains in $\mathbb{C}^{N}$ with weakly (resp. strongly) $\mathbb{C}$-extremal boundary, then so is every connected component of $D:=D_{1}\cap D_{2}$.
@@
Without loss of generality assume that $\epsilon=1$. Consider a map $\alpha:\Delta\to\mathbb{C}$ defined by $\alpha(t)=a+tu$. Since $u\neq 0$, $\alpha$ is a non-constant holomorphic map. Note that $\alpha$ is a homeomorphism from $\Delta$ onto $\alpha(\Delta)$, which is in $\partial D$. Consider the sets $S_{1}:=\{t\in\Delta:\alpha(t)\in\partial{D_{1}}\}$ and $S_{2}:=\{t\in\Delta:\alpha(t)\in\partial{D_{2}}\}$. By above decomposition, $S_{1}\cup S_{2}=\Delta$ and note that $S_{1}=\alpha^{-1}(\partial D_{1})$ and $S_{2}=\alpha^{-1}(\partial D_{2})$. Note that $\partial D_{i}\cap\partial D$ is a closed subset of $\partial D$ for $i=1,2$. By continuity of $\alpha$, $S_{1}$ and $S_{2}$ are closed sets in $\Delta$.
@@
Next, assume that $D_{1},D_{2}$ are domains in $\mathbb{C}^{N}$ with strongly $\mathbb{C}$-extremal boundary. Let $a\in\partial(D_{1}\cap D_{2})$. If possible assume that there exists $\epsilon>0$ and $u\in\mathbb{C}^{N}$ such that $a+tu\in\partial(D_{1}\cap D_{2})$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$. Note that $\partial D_{1}\cap\partial D_{2}\subset\partial(D_{1}\cap D_{2})$. Hence $\overline{D_{1}\cap D_{2}}=\overline{D_{1}}\cap\overline{D_{2}}$. If $a+tu\in\overline{D_{1}\cap D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$, then $a+tu\in\overline{D_{1}}\cap\overline{D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$. Since $D_{1}$ and $D_{2}$ are domains with strongly $\mathbb{C}$-extremal boundary, we have $u=0$. Therefore, $D_{1}\cap D_{2}$ is a domain with strongly $\mathbb{C}$-extremal boundary. ∎
@@
Thus an intersection of finitely many balls with $K$-extremal boundaries (for either case $K=\mathbb{R}$ or $\mathbb{C}$) is again a convex domain with $K$-extremal boundary. In-particular, for instance, the intersection of the $l^{1}$-ball with a suitable dilate of another $l^{p}$-ball for $p\in(1,\infty)$ can be used to get hold of a concrete instance of a domain with $\mathbb{C}$-extremal (but not smooth) boundary. As an example which is not derived out of balls with respect to the isotropic $l^{p}$-norms, we may consider the ‘anisotropic’ ball (also called a complex ellipsoid) :=$\mathcal{E}:=\{(z,w)\in\mathbb{C}^{2}\;:\;|z|^{2}+|w|^{4}<1\}$. So, there is a good variety of bounded balanced convex domains with $\mathbb{R}$-extremal and $\mathbb{C}$-extremal boundaries. The form of all holomorphic retracts of such domains passing through the point with respect to which they are balanced namely the origin, follows from the theorem due to Vesentini, already stated in the introductory section; though not in as straightforward a manner, for their products. The product $B_{1}\times B_{2}$, though bounded convex and balanced, certainly fails to satisfy the condition that its boundary points be $\mathbb{R}$-extremal, even when both factors $B_{1},B_{2}$ satisfy this condition individually. Indeed even in the simplest case when $B_{1}=B_{2}=\Delta$, the only points on the boundary of $B_{1}\times B_{2}$ which are $\mathbb{R}$-extremal are the points of its distinguished boundary, which is a rather thin subset of the boundary. As mentioned at the start, our goal is to first describe the retracts of such a product $B_{1}\times B_{2}\subset\mathbb{C}^{N_{1}+N_{2}}$ which as already noted is a bounded balanced convex domain in $\mathbb{C}^{N_{1}+N_{2}}$, indeed the unit ball with respect to the norm $\parallel(w,z)\parallel=\max\{|w|_{\mu_{1}},|z|_{{\mu_{2}}}\}$ where $\mu_{1},\mu_{2}$ denote the norms in which $B_{1},B_{2}$ are the standard (open) unit balls. As mentioned in the introduction, this theme of retracts of polyballs was initiated earlier in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, what the retracts of polyballs look like; the purpose here is to address this as systematically as we can.
@@
Thus in particular, by theorem 2.1 all retracts of the domain of our interest here, namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, retracts of our convex domain $D$ are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^3] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract. Infact, there are contractible spaces which do not admit a strong deformation retraction to any of its points, though obviously they admit a deformation retraction to each of its points. during the homotopy and for each point of time in the homotopy, the mapping is holomorphic. While this means that all retracts of $D$ are topologically trivial thereby placing strong restrictions on which complex submanifolds of $D$ can happen to be a retract, we shall see on the other hand that there is some variety as well: for instance if the factors $B_{1},B_{2}$ are either a polydisc or a Euclidean ball, there are retracts for every possible dimension which thereby have got to be holomorphically distinct. Furthermore again, not every contractible complex submanifold of $D$ passing through the origin can qualify to be a retract, for as we shall show, retracts through the origin of $B_{1}\times B_{2}$ are submanifolds in normal form in the sense that they are all graphs of holomorphic maps; as is very well-known, graphs of holomorphic functions are submanifolds which are simultaneously in parametric form as well as in level set form and the implicit function theorem allows us to put every submanifold locally in such a normal form (as graphs). In view of the foregoing theorem (and the implicit function theorem) then, retracts are obviously graphs of holomorphic maps (over some set of axes) locally; however in the specific case of retracts of $D=B_{1}\times B_{2}$ passing through the origin, we shall show that they are actually so globally i.e., each retract of the product domain $D$ is the graph of a single holomorphic map over some linear submanifold of $D$. As we shall encounter retracts in ‘graphical form’ in several cases in the sequel, we make a general definition for clarity.
@@
[…]
@@
It must be noted that the condition $\parallel df(0)\parallel=1$ essentially means that $df(0)$ is an isometry from $\mathbb{C}$ into $F$, a condition which underlies and entails linearity of certain maps to be considered later. If we drop the $\mathbb{R}$-extremal condition in the above theorem, then the conclusion about the linearity of $f$ may well fail, as shown by the simplest non-linear holomorphic map $f(z)=(z,z^{2})$ of the disc into the bidisc $\Delta^{2}$; indeed, note that $df(z)=(1,2z)$, so $df(0)=(1,0)$ thereby $\left\|df(0)\right\|_{l^{\infty}}=1$ verifying all other conditions except for the $\mathbb{R}$-extremal condition which we dropped.
The following theorem is the generalization of above Mazet theorem, which is also useful in this sequel.
@@
(Theorem 3.7 of [Mazet_Rextrm]) Let $E$ and $F$ be Banach spaces and $B_{E},B_{F}$ be the corresponding unit ball. Consider a holomorphic map $f:B_{E}\to B_{F}$ be a holomorphic map. Then we have:
@@
    If there exists $u$ of norm 1 in $E$ such that $f^{\prime}(0)\cdot u$ be $\mathbb{R}$-extremal in $\overline{B_{F}}$ then $f(0)=0$.
@@
    If, for any $u$ of norm 1 in $E,f^{\prime}(0)\cdot u$ is $\mathbb{C}$-extremal in $B_{F}$ then $f$ is a polynomial of degree at most 2.
@@
    If the two previous hypotheses are simultaneously verified, then $f$ is a linear map, $f(u)=$ $f^{\prime}(0)\cdot u$
@@
Though we shall continue making general observations for use in the sequel, we will once and for all now, specialize our ambient space from general complex manifolds or arbitrary Banach spaces to $\mathbb{C}^{N}$ equipped possibly with a norm different from the usual $l^{2}$-norm, so as to finish noting the implications of general facts for our specific case, without much digressions from this purpose. Getting to the retract $Z$ of our product domain $D=B_{1}\times B_{2}$, let $L=T_{0}Z$ which ofcourse is a complex linear subspace of $\mathbb{C}^{N_{1}+N_{2}}$; as we’ve already mentioned, we shall not be talking of trivial retracts and hence the possible dimensions of $L$ (over $\mathbb{C}$) are $11,2,\ldots,N_{1}+N_{2}-$ 1 which are also the possible dimensions of $Z$ as a complex submanifold; which of these dimensions are attained and what would be their forms then, are matters more subtle than the general features mentioned at the outset and as we shall see, actually depend on specific features of the factor domains $B_{1},B_{2}$. However, just the assumption of $\mathbb{R}$-extremality on the boundaries of $B_{1},B_{2}$ will enable us to say far more about $Z$ than just the already mentioned fact at the outset that it is a contractible submanifold of $D$. A first reason for this (as to why we shall be able to say more at all) is that we are in the setting of a complex Euclidean space which renders all the tangent spaces of $D$ to be canonically identified and for our case, parts of tangent spaces are contained within the domain $D$ itself; more precisely and in particular, there is a part of $L$ that is contained within our domain namely, the intercept $D_{L}=L\cap D$. This means that $\rho$ can be applied to this intercept (of the tangent space to $Z$) itself, which naturally gives rise to the question of the relationship between $\rho(D_{L})$ and $Z$. As to this, the version of the identity principle for complex analytic sets (as in section 5.6 of the first chapter of the book [Chirka] by Chirka, for instance) applies to ascertain that $\rho(D_{L})=Z$ where $D_{L}=L\cap D$, provided we know apriori that $\rho(D_{L})$ is a complex submanifold – obviously, $\rho(D_{L})\subset Z$ always; the role of the identity principle is to ensure the equality $\rho(D_{L})=Z$ which will be used several times in the sequel. Actually we can do better than doing this apriori verification that $\rho(D_{L})$ is a submanifold of $Z$ in each of the multiple cases that we shall be considering for the analysis of $Z$; while splitting up into multiple cases here would be seen to naturally arise owing to the difference in the forms of $Z$ that cases result in, the result $\rho(D_{L})=Z$ does not depend on any such case we may be in and can be shown to hold rather generally. As this also fits in well with our plan of first noting general facts following from from already known results, we now proceed to establish a general result about this as in theorem 3.9 below, which holds for any bounded balanced pseudoconvex domain $D$ – no $\mathbb{R}$-extremal hypotheses is required. While the proof of this will be done by borrowing an idea [Abate_Isometry], implementation of this idea requires preparatory work which itself yields another fundamental fact namely, that $D_{L}=D\cap L$ where $L=T_{0}Z$, itself is a retract of $D$, though not every linear subspace of $D$ can be its retract, unless $D$ is biholomorphic to the $l^{2}$-ball. Infact, $D_{L}$ is a linear retract of $D$ in the sense of definition 2.6.
@@
We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.7 and 2.8, allowing us to skip details here.
@@
As a corollary to this lemma, we not only get an alternative proof of corollary 2.33 – which was only for one-dimensional retracts – but also an improvement wherein there is no restriction on the dimension of the retract.
@@
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$. If there exists a point $p\in\partial D$ such that $D$ is not convex at $p$, then there does not exist a retract of $D$ passing through the origin in the direction of $p$.
@@
Assume to obtain a contradiction that there exists a retract $Z$ of $D$ passing through the origin in the direction of $p$. Let $\rho$ be any one of the retraction maps on $D$ such that ${\rm image}(\rho)=Z$. By the previous lemma 3.7, the linear projection map $\pi:=D\rho|_{0}\;:\;\mathbb{C}^{N}\to\mathbb{C}^{N}$ maps $D$ onto $D_{L}$, where $L=T_{0}Z$. Consider the affine subspace $K:=p+Ker(\pi)$ of $\mathbb{C}^{N}$. Since $D$ is not convex at $p$, there exists a point $q\in K$ such that $q\in D\cap K$. But $\pi(q)=p$, which contradicts the fact that $\pi$ maps $D$ into $D$. ∎
We return to the preceding lemma for a little clarification that saves some confusion once and for all. In case $Z$ is already a linear subspace of $D$ to begin with, the above lemma 3.7, still applies to show that a linear subspace of pseudoconvex bounded balanced domain $D$, which is apriori only known to be attainable as the image of a holomorphic (i.e., possibly non-linear) retraction map on $D$, can indeed be obtained as the image of a linear retraction map as well i.e., it is valid to say that $D_{L}$ as in the above lemma is a linear retract of $D$ without any further qualifications. We are now fully prepared to prove the highlighted general result about the equality $\rho(D_{L})=Z$ whose statement is formulated with slightly different notations than the one that we had adopted for our specific domain $D$. This is indeed the content of theorem 1.13, whose statement we lay down again here for convenience, and then prove it inspired and based on an idea in [Abate_Isometry].
@@
We first apply the foregoing lemma which realizes $D_{L}=D\cap L$ as a linear retract of $D$ by the linear projection $\pi_{L}:=D\rho(0)$. We then consider the map $F:=\pi_{L}\circ\rho_{|_{D_{L}}}$, which is a map from $D_{L}$ into itself. Note that $F$ fixes the origin and that $DF(0)$ is just the identity map as $D\rho(0)$ is the identity map on $L$ (since $L$ is a tangent space to $Z$ which is fixed pointwise by $\rho$) and $\pi_{L}$ is actually a linear retraction onto $L$ and so in particular is the identity on $L$. As $D_{L}$ is a bounded subdomain of the complex linear space $L$ containing the origin which is fixed by $F$, we may very well apply the classical uniqueness theorem of H. Cartan to deduce that $F$ itself is just the identity map. It follows that $\rho$ restricted to $D_{L}$ and $\pi_{L}$ restricted to $Z$ are inverses of each other and being holomorphic, finishes the proof of the first statement. It only remains to observe that as the inverse of $\rho_{|_{D_{L}}}$ actually happens to be the linear projection $\pi_{L}$ onto the linear subspace $D_{L}$, this means that $Z=\rho(D_{L})$ is realized essentially as the graph of a single holomorphic function over $D_{L}$ (the only linear change of coordinates that may need to be made is to orthogonalize the projection $\pi_{L}$ which may be ‘oblique’ to render $Z$ as a standard graph over $D_{L}$). ∎

### v7

@@ 2.9. Analytic and Semi-analytic sets
If $D_{1},D_{2}$ are any pair of (not necessarily convex) domains in $\mathbb{C}^{N}$ with weakly (resp. strongly) $\mathbb{C}$-extremal boundary, then so is every connected component of $D:=D_{1}\cap D_{2}$.
@@
Without loss of generality assume that $\epsilon=1$. Consider a map $\alpha:\Delta\to\mathbb{C}^{N}$ defined by $\alpha(t)=a+tu$. Since $u\neq 0$, $\alpha$ is a non-constant holomorphic map. Note that $\alpha$ is a homeomorphism from $\Delta$ onto $\alpha(\Delta)$, which is in $\partial D$. Consider the sets $S_{1}:=\{t\in\Delta:\alpha(t)\in\partial{D_{1}}\}$ and $S_{2}:=\{t\in\Delta:\alpha(t)\in\partial{D_{2}}\}$. By above decomposition, $S_{1}\cup S_{2}=\Delta$ and note that $S_{1}=\alpha^{-1}(\partial D_{1})$ and $S_{2}=\alpha^{-1}(\partial D_{2})$. Note that $\partial D_{i}\cap\partial D$ is a closed subset of $\partial D$ for $i=1,2$. By continuity of $\alpha$, $S_{1}$ and $S_{2}$ are closed sets in $\Delta$.
@@
Next, assume that $D_{1},D_{2}$ are domains in $\mathbb{C}^{N}$ with strongly $\mathbb{C}$-extremal boundary. Let $a\in\partial(D_{1}\cap D_{2})$. If possible assume that there exists $\epsilon>0$ and $u\in\mathbb{C}^{N}$ such that $a+tu\in\overline{D_{1}\cap D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$. Note that $\overline{D_{1}\cap D_{2}}=\overline{D_{1}}\cap\overline{D_{2}}$. If $a+tu\in\overline{D_{1}\cap D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$, then $a+tu\in\overline{D_{1}}\cap\overline{D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$. Since $D_{1}$ and $D_{2}$ are domains with strongly $\mathbb{C}$-extremal boundary, we have $u=0$. Therefore, $D_{1}\cap D_{2}$ is a domain with strongly $\mathbb{C}$-extremal boundary. ∎
@@
Thus an intersection of finitely many balls with $K$-extremal boundaries (for either case $K=\mathbb{R}$ or $\mathbb{C}$) is again a convex domain with $K$-extremal boundary. In-particular, for instance, the intersection of the $l^{1}$-ball with a suitable dilate of another $l^{p}$-ball for $p\in(1,\infty)$ can be used to get hold of a concrete instance of a domain with $\mathbb{C}$-extremal (but not smooth) boundary. As an example which is not derived out of balls with respect to the $l^{p}$-norms (which are symmetric under a permutation of coordinates, despite not being strictly isotropic), we may for example, consider the ‘more’ anisotropic ball $\mathcal{E}:=\{(z,w)\in\mathbb{C}^{2}\;:\;|z|^{2}+|w|^{4}<1\}$ referred to as a ‘complex ellipsoid’ or an ‘egg’ domain . So, there is a good variety of bounded balanced convex domains with $\mathbb{R}$-extremal and $\mathbb{C}$-extremal boundaries. The form of all holomorphic retracts of such domains passing through the point with respect to which they are balanced namely the origin, follows from the theorem due to Vesentini, already stated in the introductory section; though not in as straightforward a manner, for their products. The product $B_{1}\times B_{2}$, though bounded convex and balanced, certainly fails to satisfy the condition that its boundary points be $\mathbb{R}$-extremal, even when both factors $B_{1},B_{2}$ satisfy this condition individually. Indeed even in the simplest case when $B_{1}=B_{2}=\Delta$, the only points on the boundary of $B_{1}\times B_{2}$ which are $\mathbb{R}$-extremal are the points of its distinguished boundary, which is a rather thin subset of the boundary. As mentioned at the start, one of our goals is to first describe the retracts of such a product $B_{1}\times B_{2}\subset\mathbb{C}^{N_{1}+N_{2}}$, the unit ball with respect to the norm $\parallel(w,z)\parallel=\max\{|w|_{\mu_{1}},|z|_{{\mu_{2}}}\}$ where $\mu_{1},\mu_{2}$ denote the norms in which $B_{1},B_{2}$ are the standard (open) unit balls.
@@
Thus in particular, by theorem 2.1 all retracts of the domains as in theorem 1.17 namely $D:=B_{1}\times B_{2}$, are submanifolds thereof and moreover are all contractible since $B_{1}\times B_{2}$ is; as already indicated, the retracts in the case $B_{1},B_{2}$ are convex, are also strong holomorphic deformation retracts – indeed, the linear homotopy between the identity map of $D$ and any retraction gives a homotopy in which the retract remains pointwise fixed  [^9] In general, while every (continuous) retract of a contractible topological space $X$ is a deformation retract much as any self-map of $X$ is homotopic to the identity, not every retract is a strong deformation retract. Infact, there are contractible spaces which do not admit a strong deformation retraction to any of its points, though obviously they admit a deformation retraction to each of its points. during the homotopy and for each point of time in the homotopy, the mapping is holomorphic. While this means that all retracts of $D$ are topologically trivial thereby placing strong restrictions on which complex submanifolds of $D$ can happen to be a retract, we shall see on the other hand that there is some variety as well: for instance if the factors $B_{1},B_{2}$ are either a polydisc or a Euclidean ball, there are retracts for every possible dimension which thereby have got to be holomorphically distinct. Furthermore again, not every contractible complex submanifold of $D$ passing through the origin can qualify to be a retract, for as we shall show, retracts through the origin of $B_{1}\times B_{2}$ are submanifolds in normal form in the sense that they are all graphs of holomorphic maps; as is very well-known, graphs of holomorphic functions are submanifolds which are simultaneously in parametric form as well as in level set form and the implicit function theorem allows us to put every submanifold locally in such a normal form (as graphs). In view of the foregoing theorem (and the implicit function theorem) then, retracts are obviously graphs of holomorphic maps (over some set of axes) locally; however in the specific case of retracts of $D=B_{1}\times B_{2}$ passing through the origin, we shall show that they are actually so globally i.e., each retract of the product domain $D$ is the graph of a single holomorphic map over some linear submanifold of $D$. As we shall encounter retracts in ‘graphical form’ in several cases in the sequel, we make a general definition for clarity.
@@
Before delving into a detailed analysis of retracts through the origin of balanced domains $D$, one may like to see at this stage what can be said if at all, about such retracts of bounded balanced convex domains $D$ in general. Let us begin with the trivia that retractions $\rho$ associated to such retracts of $D$ fix the origin $\rho(0)=0$, and are self-maps of $D$; note also that $D$ being a bounded balanced convex domain, is the unit ball with respect to some norm $\left\|\cdot\right\|$ on $\mathbb{C}^{N}$, indeed its Minkowski functional. This already prompts us to apply an appropriate version of the Schwarz lemma on $\rho$. A befitting version for this context alongwith an elementary proof is provided by proposition C of the article [Simh] by Simha, by which we get $\left\|D\rho(0)\right\|\leq 1$. However, unlike the classical Schwarz lemma, equality herein need not always imply linearity of the map, to attain which, additional conditions need to be imposed. This brings us to the version of the Schwarz lemma due to Mazet mentioned at the outset, which we shall generalize below so as to make recurrent use of it in the sequel.
@@
It must be noted that the condition $\parallel df(0)\parallel=1$ essentially means that $df(0)$ is an isometry from $\mathbb{C}$ into $F$, a condition which underlies and entails linearity of certain maps to be considered later. If we drop the $\mathbb{R}$-extremal condition in the above theorem, then the conclusion about the linearity of $f$ may well fail, as shown by the simplest non-linear holomorphic map $f(z)=(z,z^{2})$ of the disc into the bidisc $\Delta^{2}$; indeed, note that $df(z)=(1,2z)$, so $df(0)=(1,0)$ thereby $\left\|df(0)\right\|_{l^{\infty}}=1$ verifying all other conditions except for the $\mathbb{R}$-extremal condition which we dropped. The following extension also due to Mazet, which completely drops restrictions on the dimension of the source of the map is no less relevant in the sequel. We shall generalize this to some extent below for our needs.
@@
(Theorem 3.7 of [Mazet_Rextrm]) Let $E,F$ be Banach spaces with their unit balls denoted $B_{E},B_{F}$ respectively. Let $f:B_{E}\to B_{F}$ be any holomorphic map. Then we have the following:
@@
    If there exists $u$ of norm 1 in $E$ such that $f^{\prime}(0)\cdot u$ is $\mathbb{R}$-extremal in $\overline{B_{F}}$ then $f(0)=0$.
@@
    If, for every $u$ of norm $1$ in $E$, $f^{\prime}(0)\cdot u$ is $\mathbb{C}$-extremal in $B_{F}$ then $f$ is a polynomial of degree at most 2.
@@
    If the two previous hypotheses are simultaneously satisfied, then $f$ is a linear map, $f(u)=$ $f^{\prime}(0)\cdot u$.
An analogue of the above for the non-convex case can be derived from some standard ideas from chapter $11$ of [Jrncki_invrnt_dst], which we do now. This will quickly lead to a generalization of Vesentini’s theorem for the non-convex case, which will however be derived a bit later after theorem 3.10 where it is more relevant.
Theorem 3.6.
Let $D$ be a balanced pseudoconvex (not necessarily bounded) domain in $\mathbb{C}^{N}$ and $\varphi:\Delta\to D$ be a holomorphic map. If $\varphi(0)=0$ and $\varphi^{\prime}(0)$ happens to be a holomorphic extreme boundary point in $D$, then $\varphi$ is a linear map: $\varphi(t)=\varphi^{\prime}(0)t$. More generally, if $G$ is a bounded balanced pseudoconvex domain in $\mathbb{C}^{M}$ and $F:G\to D$ be a holomorphic map with the property that $F(0)=0$ and for all $u\in\partial G$, $DF_{|_{0}}(u)$ is a holomorphic extreme boundary point of $D$, then $F$ is linear.
Proof.
We start with the holomorphic map $\varphi:\Delta\to D$ as in the first part of the statement i.e., satisfying $\varphi(0)=0$ with $\varphi^{\prime}(0)=p$ being a holomorphic extreme boundary point of $D$. As $\varphi(0)=0$, we can write $\varphi(t)=t\tilde{\varphi}(t)$, where $\tilde{\varphi}\in\mathcal{O}(\Delta,\mathbb{C}^{N})$. Since $\varphi$ maps $\Delta$ into $D$, we have
$h(\varphi(t))=|t|h\circ\tilde{\varphi}(t)<1.$
Because the above inequality holds for all $t\in\Delta$, we obtain a similar inequality (but with the possibility of equality): $h\circ\tilde{\varphi}(t)\leq 1$ for all $t\in\Delta$. Note that
$\tilde{\varphi}(0)=\varphi^{\prime}(0)=p.$
This implies that $h\circ\tilde{\varphi}(0)=h(p)=1$. By the maximum principle applied to the subharmonic function $h\circ\tilde{\varphi}$, we get that $h\circ\tilde{\varphi}\equiv 1$. This shows that ${\rm image}(\tilde{\varphi})\subset\partial D$. By the hypothesis that $p\in\partial D$ is a holomorphic extreme boundary point of $D$, we obtain that $\tilde{\varphi}\equiv p$ and hence we conclude that $\varphi(t)=tp$ for all $t\in\Delta$.
We shall now use the above to prove the general case expressed in-terms of the map $F$ as in the statement of the theorem. Let $p$ be an arbitrary point in $\partial G$. Consider the holomorphic map $\varphi:\Delta\to D$ defined by $\varphi(t)=F(tp)$. Note that $\varphi(0)=0$ and $\varphi^{\prime}(0)=DF_{|_{0}}(p)$ is a holomorphic extreme boundary point of $D$. By the special case $G=\Delta$ already dealt with, we obtain that $\varphi(t)=tDF_{|_{0}}(p)$. As $p$ was arbitrarily chosen, this means that $F(tp)=DF_{|_{0}}(tp)$ for all $p$ in the boundary of the bounded balanced domain $G$ (and all $t\in\Delta$). It follows thereby that $F(z)=DF_{|_{0}}(z)$, finishing the proof that $F$ is a linear map. ∎
@@
We now get to the retract $Z$ of our product domain $D=B_{1}\times B_{2}$ and denote its tangent space at the origin $T_{0}Z$ by $L$, which ofcourse is a complex linear subspace of $\mathbb{C}^{N_{1}+N_{2}}$; as we have already mentioned, we shall not be talking of trivial retracts and hence the possible dimensions of $L$ (over $\mathbb{C}$) are $1,2,\ldots,N_{1}+N_{2}-1$ which are also the possible dimensions of $Z$ as a complex submanifold. Which of these dimensions are attained and what would be their forms then, are matters more subtle than the general features mentioned at the outset and as we shall see, actually depend on specific features of the factor domains $B_{1},B_{2}$. However, just the assumption of $\mathbb{R}$-extremality on the boundaries of $B_{1},B_{2}$ will enable us to say far more about $Z$ than just the already mentioned fact at the outset that it is a contractible submanifold of $D$. A first reason for this (as to why we shall be able to say more at all) is that we are in the setting of a complex Euclidean space which renders all the tangent spaces of $D$ to be canonically identified and for our case, parts of tangent spaces are contained within the domain $D$ itself; more precisely and in particular, there is a part of $L$ that is contained within our domain namely, the intercept $D_{L}=L\cap D$. This means that $\rho$ can be applied to this intercept (of the tangent space to $Z$) itself, which naturally gives rise to the question of the relationship between $\rho(D_{L})$ and $Z$. As to this, the version of the identity principle for complex analytic sets (as in section 5.6 of the first chapter of the book [Chirka] by Chirka, for instance) applies to ascertain that $\rho(D_{L})=Z$ where $D_{L}=L\cap D$, provided we know apriori that $\rho(D_{L})$ is a complex submanifold – obviously, $\rho(D_{L})\subset Z$ always; the role of the identity principle is to ensure the equality $\rho(D_{L})=Z$ which will be used several times in the sequel. Actually we can do better than doing this apriori verification that $\rho(D_{L})$ is a submanifold of $Z$ in each of the multiple cases that we shall be considering for the analysis of $Z$; while splitting up into multiple cases here would be seen to naturally arise owing to the difference in the forms of $Z$ that cases result in, the result $\rho(D_{L})=Z$ does not depend on any such case we may be in and can be shown to hold rather generally. As this also fits in well with our plan of first noting general facts following from already known results, we now proceed to establish a general result about this as in theorem 3.10 below, which holds for any bounded balanced pseudoconvex domain $D$ – no $\mathbb{R}$-extremal hypotheses is required. While the proof of this will be done by borrowing an idea from [Abate_Isometry], implementation of this idea requires preparatory work which itself yields another fundamental fact namely, that $D_{L}=D\cap L$ where $L=T_{0}Z$, itself is a retract of $D$, though not every linear subspace of $D$ can be its retract, unless $D$ is biholomorphic to the $l^{2}$-ball. Indeed, $D_{L}$ is a linear retract of $D$ in the sense of definition 2.10.
@@
We can now progress towards our aforementioned general result on fully attaining $Z$ as the image of $L$ under $\rho$, the proof of which requires the subsequent lemmas which in turn follow fairly directly from the ideas in the pair of remarks 2.11 and 2.12, allowing us to skip details here.
@@
As a corollary to this lemma, we give a proof of corollary 2.38 not because this proof avoids geodesics but because it brings out the reason directly.
@@
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$. If there exists a point $p\in\partial D$ such that $D$ is not convex at $p$, then there does not exist any one-dimensional retract of $D$ passing through the origin in the direction of $p$.
@@
Assume to obtain a contradiction that there exists a one-dimensional retract $Z$ of $D$ passing through the origin in the direction of $p$. Let $\rho$ be any one of the retraction maps on $D$ such that ${\rm image}(\rho)=Z$. By the previous lemma 3.8, the linear projection map $\pi:=D\rho_{|_{0}}\;:\;\mathbb{C}^{N}\to\mathbb{C}^{N}$ maps $D$ onto $D_{L}$, where $L=T_{0}Z$. Consider the affine subspace $K:=p+\ker(\pi)$ of $\mathbb{C}^{N}$. Note that $K$ is a complex hyperplane at $p$. Since $D$ is not convex at $p$, there exists a point $q\in K$ such that $q\in D\cap K$. But $\pi(q)=p$, which contradicts the fact that $\pi$ maps $D$ into $D$. ∎
We return to the preceding lemma for a little clarification that saves some confusion once and for all. In case $Z$ is already a linear subspace of $D$ to begin with, the above lemma 3.8, still applies to show that a linear subspace of pseudoconvex bounded balanced domain $D$, which is apriori only known to be attainable as the image of a holomorphic (i.e., possibly non-linear) retraction map on $D$, can indeed be obtained as the image of a linear retraction map as well i.e., it is valid to say that $D_{L}$ as in the above lemma is a linear retract of $D$ without any further qualifications. We are now fully prepared to prove the highlighted general result about the equality $\rho(D_{L})=Z$ whose statement is formulated with slightly different notations than the one that we had adopted for our specific domain $D$. This is indeed the content of theorem 1.15, whose statement we lay down again here for convenience, and then prove it based on an idea in [Abate_Isometry].
@@
We first apply the foregoing lemma which realizes $D_{L}=D\cap L$ as a linear retract of $D$ by the linear projection $\pi_{L}:=D\rho(0)$. We then consider the map $F:=\pi_{L}\circ\rho_{|_{D_{L}}}$, which is a map from $D_{L}$ into itself. Note that $F$ fixes the origin and that $DF(0)$ is just the identity map as $D\rho(0)$ is the identity map on $L$ (since $L$ is a tangent space to $Z$ which is fixed pointwise by $\rho$) and $\pi_{L}$ is actually a linear retraction onto $L$ and so in particular is the identity on $L$. As $D_{L}$ is a bounded subdomain of the complex linear space $L$ containing the origin which is fixed by $F$, we may very well apply the classical uniqueness theorem of H. Cartan to deduce that $F$ itself is just the identity map. It follows that $\rho$ restricted to $D_{L}$ and $\pi_{L}$ restricted to $Z$ are inverses of each other and being holomorphic, finishes the proof of the first statement. It only remains to observe that as the inverse of $\rho_{|_{D_{L}}}$ actually happens to be the linear projection $\pi_{L}$ onto the linear subspace $D_{L}$ along its kernel $M=\ker\left(D\rho(0)\right)$, this means that $Z=\rho(D_{L})$ is realized essentially as the graph of a (single) holomorphic map over $D_{L}$ taking values in $M=\ker\left(D\rho(0)\right)$. ∎